# Constraints from Entanglement Wedge Nesting for Holography at a Finite Cutoff

### v2

@@ 4.2.2 Connected Extremal Surfaces
To find the connected extremal surfaces we extremize the area function in Kruskal coordinates,
@@
When we do our analysis of entanglement wedge nesting we will find it useful to work in Schwarzschild coordinates. To find the expressions for the connected RT curves we start with Eq. (75) and make use of the coordinate transformation that takes exterior regions in Kruskal coordinates to Schwarzschild coordinates given by Eq. (58). The result is the following expression (see Appendix C.2 for details)
@@ 4.3 Construction of Entanglement Wedges
We now turn to discussing the entanglement wedges associated to the regions $A_{\text{Bdry}}$, $A_{\text{Br}}$, and $A_{\text{Bdry}}\cup A_{\text{Br}}$. Starting with $A_{\text{Bdry}}$, we use $\Sigma_{A_{\text{Bdry}}}$ to denote a partial Cauchy surface with boundary $\partial\Sigma_{A_{\text{Bdry}}}=A_{\text{Bdry}}\cup\chi_{\text{ther}}(A_{\text{Bdry}})$. We define the entanglement wedge $\mathcal{W}_{E}(A_{\text{Bdry}})$ of $A_{\text{Bdry}}$ as the domain of dependence of $\Sigma_{A_{\text{Bdry}}}$. A convenient prescription which can be used to explicitly construct $\mathcal{W}_{E}(A_{\text{Bdry}})$ is by considering the a family of null geodesics beginning on $\chi_{\text{ther}}(A_{\text{Bdry}})$ which are also orthogonal to $\chi_{\text{ther}}(A_{\text{Bdry}})$ and evolve them towards the boundary. The set of these null geodesics enclose a co-dimension 0 region which is exactly $\mathcal{W}_{E}(A_{\text{Bdry}})$ as shown in Figure 13(a).
@@
4.4 Conditions for EWN
In this section we will derive constraints on the parameters defining the separation of boundary and brane subregions in Kruskal time such that $\mathcal{W}_{E}(A_{\text{Br}})\subset\mathcal{W}_{E}(A_{\text{Br}}\cup A_{\text{Bdry}})$ in the connected phase.
Just as in the Poincare AdS3 analysis in Section 3.1 a sufficient condition for EWN is given by requiring that all null geodesics originating from $\chi_{\text{con}}(A_{\text{Bdry}}\cup A_{\text{Br}})$ are unable to reach $\mathcal{W}_{E}(A_{\text{Br}})$, i.e. no causal signals originating from $\chi_{\text{con}}(A_{\text{Bdry}}\cup A_{\text{Br}})$ can influence events in $\mathcal{W}_{E}(A_{\text{Br}})$. Due to the trivial trajectory of the connected surface in the $xz$-plane, it suffices to ask when the points on the connected surface are spacelike separated from $\mathcal{W}_{E}(\text{Vir}(A_{\text{Br}}))$. The reason is that in our particular setup of brane and boundary intervals the connected surface always lies entirely the the $\mathcal{R}_{M}$ region (depicted in Figure 20).
@@ 4.4 Conditions for EWN
and solutions are [^21] Details of getting solution given in Appendix C.5.,
@@
The tightest interval occurs when we take the limit of the expressions above as $\Xi\to 1$ (i.e. when $a\to\infty$) to give [^22] In Appendix C.6 we show that the connected phase will dominate in this regime. Intuitively, the thermal extremal surface area goes to infinity while the connected surface area remains finite as $a\to\infty$.,
@@
4.5 Constraints from EWN
Now that we have derived a sufficient condition for when EWN is respected in the connected phase we can continue and describe a nice geometric interpretation of what the tightest constraints given in Eq. (84) mean. First lets recall that a priori we just required that $|c_{t}|<1$ so the connected RT surfaces are spacelike. Now suppose we have a connected curve which starts on the left boundary at some Kruskal time $|\tau_{\text{Bdry}}|<\pi/2$. For $c_{t}=0$ the connected surface in Kruskal coordinates has a trivial horizontal profile and will end on the brane in the right exterior at the same Kruskal time as it started with at the boundary. However for more general values of $c_{t}$ less than unity one can imagine that there are certain values of $c_{t}$ where the surface would go through the past or future singularity before it hits the brane.
@@ 4.5 Constraints from EWN
It turns out the the condition that EWN hold for intervals on the brane in the right exterior exactly excludes the configurations that would allow the connected surface to go through the singularity. In particular, if we fix the location of a constant time interval in the right exterior on the brane (i.e. fix $\tau_{\text{Br}}$ and $y_{\text{Br}}$) then the condition that EWN holds translates to
@@ 5 Concluding Remarks
Our understanding of holography at a finite cutoff is poor and we hope that a careful study of properties of entanglement wedges in the presence of a cutoff can shed light on the properties of holographic duals to finite subregions. It would in particular be interesting to generalize our results to more complicated geometries and interval configurations, also in higher dimensions. Optimistically, this could provide a set of rules which determine under which conditions cutoff and boundary subregions can be treated as truly independent. It is interesting to note that violations of the conditions we obtained for EWN throughout this paper have always started in arbitrarily small bulk neighborhoods of the entangling surfaces. Perhaps this will always be the case if a non-trivial violation of EWN occurs, such that it could be sufficient to focus on those regions which would significantly simplify the analyses. Even if this is not generally the case one might at least formulate some concrete necessary conditions for EWN to hold again, this may provide insight as to how one can naturally identify independent subregions in cutoff holography. It may also be interesting to explore is how much our conclusions are affected through the particular model of cutoff surface we choose. Here we chose surfaces that are often employed in doubly holography setups and one might wonder if things may change for more general choices of cutoff surfaces [^23] It was pointed out in Omiya:2021olc that the ETW branes considered in double holography fails to satisfy the Gao-Wald theorem Gao:2000ga . We believe that our results can be regarded as a consequence of such violations of the Gao-Wald theorem. Based on this, we might speculate that more general cutoff surfaces also feature such violations of the Gao-Wald theorem and therefore our conclusions may carry over in a similar manner in more general contexts..
@@
Finally, another interesting avenue of research goes back to our discussion near the end of Section 1.1 of boundary conditions on the brane. In particular, when we impose Neumann boundary conditions on the brane we should think of the constraints that we obtain from EWN as constraints on brane world gravity. The rough procedure would be to understand how the location of the brane subregion, $B$, [^24] We fix the location of $A$ on a conformal boundary and location of $B$ is determined dynamically by understanding where the connected RT surface ends on the brane which should depend on the gravitational couplings on the brane. depends on various gravitational couplings in the brane-world theory of gravity. By precisely understanding this relation one could then straightforwardly translate the constraints we found here to constraints on gravity couplings. We leave such an investigation for later work.
@@ A.2 Spacelike Separated Points From Line Segments AdS3
To understand the the set of points that are causally disconnected from the line segment we can consider an arbitrary point on the line segment $(t(\lambda_{0}),x(\lambda_{0}),z(\lambda_{0}))$, where $\lambda_{0}\in(\lambda_{L},\lambda_{R})$. We consider the lightcone centered at the point
@@
where the $\pm$ defines the two branches of the RT surface when when parameterize in terms of $z$. We can plug this into $t(x)$ to get
@@
where $k=\frac{x_{+}-x_{-}}{t_{+}-t_{-}}$ and $x\in\left[\frac{x_{+}+x_{-}}{2}-\frac{t_{+}-t_{-}}{2},\frac{x_{+}+x_{-}}{2}+\frac{t_{+}-t_{-}}{2}\right]$. This allows us to express the set of spacelike separated points from the extremal curve towards the AdS3 boundary [^25] In Appendix A.2 we show that to understand the points that are spacelike separated to a line segment we can consider null surfaces generated by emitting null geodesics orthogonally from the RT curve. as
@@
where $\Delta t=t_{b}-t_{a}$, $\Delta x_{i}=a_{i}+b_{i}\sin\theta_{0}$, and $\Delta z_{i}=b_{i}\cos\theta_{0}$. The $\pm$ signs appearing in Eq. (122) determine if the null points are have larger $z$ than $\chi_{i}$ (given by “$+$” sign) or a smaller $z$ than $\chi_{i}$ (given by “--” sign) , in the $xz$-plane for each fixed time slice. What we plot in the Figure 6 is $z_{\text{Null},i}(x)=\left[z^{(+)}_{\text{Null},i}(x,t=T)\cup z^{(-)}_{\text{Null},i}(x,t=T)\right]$ at various fixed time slices $t=T$ [^26] NOTE: The Heaviside step function, $\Theta$, used in Eq. (122) has a value of $1/2$ when the argument is zero (i.e. $\Theta(0)=\frac{1}{2}$).
@@ B.1 Analysis of Conditions for EWN
where $z_{i}(x)$ and $t_{i}(x)$ are given in Eq. (12). Along $\chi_{i}$ we have have the following relation between $x$ and $|t_{i}(x)-t_{a}|$,
@@
Next we consider the conditions given in Eq. (35). For $\chi_{i}$ we we need to consider,
@@
$\begin{split}&\left[|\delta t_{i}|+R_{b}\right]^{2}-\left[b_{i}\sin\theta_{0}-\frac{|\delta t_{i}|}{|k_{i}|}-\frac{b_{1}+b_{2}}{2\sin\theta_{0}}\right]^{2}-z_{1}\left(b_{i}\sin\theta_{0}-\frac{|\delta t_{i}|}{|k_{i}|}\right)^{2}\\ &=\frac{|\delta t_{i}|(a_{i}+b_{i}\sin\theta_{0})}{|k_{i}|}\left[\left[|k_{i}|+\frac{R_{b}}{a_{i}+b_{i}\sin\theta_{0}}\right]^{2}-\frac{\left(2a_{i}+\frac{b_{1}+b_{2}}{\sin\theta_{0}}\right)^{2}}{4(a_{i}+b_{i}\sin\theta_{0})^{2}}\right]<0.\\ \end{split}$ (129)
@@
The set of of terms in in the second line are zero due to Eq. (133). We arrive at the following conclusion,
@@
So whenever this holds we will know that at least in a neighborhood of the point where the brane and $\chi_{1}$ intersect $\chi_{1}\cap\mathcal{R}_{M}=\varnothing$ [^27] The reader may be concerned that the expression for $k_{1,*}$ is not always going to be real. This is indeed true so the reader should really interpret the inequality is being true when $k_{1,*}\in\mathbb{R}$. In the case when it is not real we can still show that $\chi_{1}\cap\mathcal{R}_{M}=\varnothing$..
@@
The reason for this is because of the the result we wrote in Eq. (143), it shows that $-z_{1}(x)|_{k_{1}=0}\geq-z_{1}(x)\Rightarrow z_{\sigma_{L}}(x)-z_{1}(x)|_{k_{1}=0}\geq z_{\sigma_{L}}(x)-z_{1}(x)$. Therefore, satisfying the bounds written in Eq. (147) for $k_{1}=0$ is sufficient to have a bound for general values of $k_{1}$. A simple computation reveals that
@@
    • $|k_{1}|\geq\text{Re}\left[k_{1,*}\right]$ [^30] Here we are combining the cases when $k_{1,*}$ is real or imaginary by simply putting bounds using the real part of $k_{1,*}$. .
@@ B.3 Condition for Dominance of Connected Phase
For example, in Figure 16 we plot both $\Delta t_{A,EWN}^{2}-\Delta t_{\text{con}}^{2}$ (left frame) and $\Delta t_{A,EWN}^{2}-\Delta t_{\text{con}}^{2}$ (right frame) as functions of $\theta_{0}$ for fixed $b_{1,2}$ and $a_{+}$, and each colored line in the graph corresponds to a particular choice of $a_{-}$ [^31] The reader might be wondering why we take the difference of squares rather than difference of the quantities themselves. It is just to reduce the number of square roots in our expression which can cause reduce numerical error when we get close to machine precision. We know the quantities themselves are always non-negative so can do this with. ut affecting our conclusions.. At first glance, we can see for certain choices of parameters the lines can be above or below the $x$-axis. We can also see that when comparing between the left and right frames, lines of the same color (i.e. the lines that are plotted for same parameters $b_{1,2}$ and $a_{\pm}$) are above or below the $x$-axis for the exact same interval(s) in $\theta_{0}$. This can be seen visually in the plots, and verified through more precise numerical calculations. This is not just true for the particular sample we show but also more generally. This in fact is likely not a coincidence; it is a consequence of another fact that we numerically verified for many different choices of parameters. Which is, the roots of $\Delta t_{A,EWN}-\Delta t_{\text{con}}$ and $\Delta t_{2,EWN}-\Delta t_{\text{con}}$ are precisely given by solutions to $2a_{-}=\ell_{*}$, where we introduced $\ell_{*}$ in Eq. (48). In fact, recalling that by definition $\ell_{*}=2a_{-}\Leftrightarrow\Delta t_{A,EWN}=\Delta t_{2,EWN}$ makes it very clear why the roots match for the same colored lines in Figure 16.

### v3

@@ 4.2.2 Connected Extremal Surfaces
To find the connected extremal surfaces we extremize the area functional in Kruskal coordinates,
@@
When we do our analysis of entanglement wedge nesting we will find it useful to work in Schwarzschild coordinates. To find the expressions for the connected extremal curves we start with Eq. (75) and make use of the coordinate transformation that takes exterior regions in Kruskal coordinates to Schwarzschild coordinates given by Eq. (58). The result is the following expression (see Appendix C.2 for details)
@@ 4.3 Construction of Entanglement Wedges
We now turn to discussing the entanglement wedges associated to the regions $A_{\text{Bdry}}$, $A_{\text{Br}}$, and $A_{\text{Bdry}}\cup A_{\text{Br}}$. Starting with $A_{\text{Bdry}}$, we use $\Sigma_{A_{\text{Bdry}}}$ to denote a partial Cauchy surface with boundary $\partial\Sigma_{A_{\text{Bdry}}}=A_{\text{Bdry}}\cup\chi_{\text{ther}}(A_{\text{Bdry}})$. We define the entanglement wedge $\mathcal{W}_{E}(A_{\text{Bdry}})$ of $A_{\text{Bdry}}$ as the domain of dependence of $\Sigma_{A_{\text{Bdry}}}$. A convenient prescription which can be used to explicitly construct $\mathcal{W}_{E}(A_{\text{Bdry}})$ involves the family of null geodesics beginning on $\chi_{\text{ther}}(A_{\text{Bdry}})$ which are also orthogonal to $\chi_{\text{ther}}(A_{\text{Bdry}})$ and evolve them towards the boundary. The set of these null geodesics enclose a co-dimension 0 region which is exactly $\mathcal{W}_{E}(A_{\text{Bdry}})$ as shown in Figure 13(a).
@@
4.4 Condition for EWN
In this section we will derive constraints on the parameters defining the separation of boundary and brane subregions in Kruskal time such that $\mathcal{W}_{E}(A_{\text{Br}})\subset\mathcal{W}_{E}(A_{\text{Br}}\cup A_{\text{Bdry}})$ in the connected phase and also provide the geometric interpretation of the constraints.
Just as in the Poincare AdS3 analysis in Section 3.1 a sufficient condition for EWN is given by requiring that all null geodesics originating from $\chi_{\text{con}}(A_{\text{Bdry}}\cup A_{\text{Br}})$ are unable to reach $\mathcal{W}_{E}(A_{\text{Br}})$, i.e. no causal signals originating from $\chi_{\text{con}}(A_{\text{Bdry}}\cup A_{\text{Br}})$ can influence events in $\mathcal{W}_{E}(A_{\text{Br}})$. Due to the trivial trajectory of the connected surface in the $xz$-plane, it suffices to ask when the points on the connected surface are spacelike separated from $\mathcal{W}_{E}(\text{Vir}(A_{\text{Br}}))$. The reason is that in our particular setup of brane and boundary intervals the connected surface always lies entirely the $\mathcal{R}_{M}$ region [^24] Since we are dealing with all the surfaces lying in $\mathcal{R}_{M}$ the sufficient condition we derive will also be necessary analogous to the AdS3 case. (depicted in Figure 20).
@@ 4.4 Condition for EWN
and solutions are, [^25] Details of getting solution given in Appendix C.5.
@@
The tightest interval occurs when we take the limit of the expressions above as $\Xi\to 1$ (i.e. when $a\to\infty$) to give, [^26] In Appendix C.6 we show that the connected phase will dominate in this regime. Intuitively, the thermal extremal surface area goes to infinity while the connected surface area remains finite as $a\to\infty$.
@@
Now that we have derived a condition for when EWN is respected in the connected phase we can continue and describe a nice geometric interpretation of what the tightest constraints given in Eq. (84) mean. First lets recall that a priori we just required that $|c_{t}|<1$ so the connected RT surfaces are spacelike. Now suppose we have a connected curve which starts on the left boundary at some Kruskal time $|\tau_{\text{Bdry}}|<\pi/2$. For $c_{t}=0$ the connected surface in Kruskal coordinates has a trivial horizontal profile and will end on the brane in the right exterior at the same Kruskal time as it started with at the boundary. However for more general values of $c_{t}$ less than unity one can imagine that there are certain values of $c_{t}$ where the surface would go through the past or future singularity before it hits the brane.
@@
It turns out the condition that EWN hold for intervals on the brane in the right exterior exactly excludes the configurations that would allow the connected surface to go through the singularity. In particular, if we fix the location of a constant time interval in the right exterior on the brane (i.e. fix $\tau_{\text{Br}}$ and $y_{\text{Br}}$) then the condition that EWN holds translates to
@@ 5 Concluding Remarks
Our understanding of holography at a finite cutoff is poor and we hope that a careful study of properties of entanglement wedges in the presence of a cutoff can shed light on the properties of holographic duals to finite subregions. It would in particular be interesting to generalize our results to more complicated geometries and interval configurations, also in higher dimensions. Optimistically, this could provide a set of rules which determine under which conditions cutoff and boundary subregions can be treated as truly independent. It is interesting to note that violations of the conditions we obtained for EWN throughout this paper have always started in arbitrarily small bulk neighborhoods of the entangling surfaces. Perhaps this will always be the case if a non-trivial violation of EWN occurs, such that it could be sufficient to focus on those regions which would significantly simplify the analyses. Even if this is not generally the case one might at least formulate some concrete necessary conditions for EWN to hold again, this may provide insight as to how one can naturally identify independent subregions in cutoff holography. It may also be interesting to explore is how much our conclusions are affected through the particular model of cutoff surface we choose. Here we chose surfaces that are often employed in doubly holography setups and one might wonder if things may change for more general choices of cutoff surfaces. [^27] It was pointed out in Omiya:2021olc that the ETW branes considered in double holography fails to satisfy the Gao-Wald theorem Gao:2000ga . We believe that our results can be regarded as a consequence of such violations of the Gao-Wald theorem. Based on this, we might speculate that more general cutoff surfaces also feature such violations of the Gao-Wald theorem and therefore our conclusions may carry over in a similar manner in more general contexts.
@@
Finally, another interesting avenue of research goes back to our discussion near the end of Section 1.1 of boundary conditions on the brane. In particular, when we impose Neumann boundary conditions on the brane we should think of the constraints that we obtain from EWN as constraints on brane world gravity. The rough procedure would be to understand how the location of the brane subregion, $B$, depends on various gravitational couplings in the brane-world theory of gravity. [^28] We fix the location of $A$ on a conformal boundary and location of $B$ is determined dynamically by understanding where the connected RT surface ends on the brane which should depend on the gravitational couplings on the brane. By precisely understanding this relation one could then straightforwardly translate the constraints we found here to constraints on gravity couplings. We leave such an investigation for later work.
@@ A.2 Spacelike Separated Points From Line Segments AdS3
To understand the set of points that are causally disconnected from the line segment we can consider an arbitrary point on the line segment $(t(\lambda_{0}),x(\lambda_{0}),z(\lambda_{0}))$, where $\lambda_{0}\in(\lambda_{L},\lambda_{R})$. We consider the lightcone centered at the point
@@
where the $\pm$ defines the two branches of the RT surface when parameterize in terms of $z$. We can plug this into $t(x)$ to get
@@
where $k=\frac{x_{+}-x_{-}}{t_{+}-t_{-}}$ and $x\in\left[\frac{x_{+}+x_{-}}{2}-\frac{t_{+}-t_{-}}{2},\frac{x_{+}+x_{-}}{2}+\frac{t_{+}-t_{-}}{2}\right]$. This allows us to express the set of spacelike separated points from the extremal curve towards the AdS3 boundary as [^29] In Appendix A.2 we show that to understand the points that are spacelike separated to a line segment we can consider null surfaces generated by emitting null geodesics orthogonally from the RT curve.
@@
where $\Delta t=t_{b}-t_{a}$, $\Delta x_{i}=a_{i}+b_{i}\sin\theta_{0}$, and $\Delta z_{i}=b_{i}\cos\theta_{0}$. The $\pm$ signs appearing in Eq. (122) determine if the null points are have larger $z$ than $\chi_{i}$ (given by “$+$” sign) or a smaller $z$ than $\chi_{i}$ (given by “--” sign) , in the $xz$-plane for each fixed time slice. What we plot in the Figure 6 is $z_{\text{Null},i}(x)=\left[z^{(+)}_{\text{Null},i}(x,t=T)\cup z^{(-)}_{\text{Null},i}(x,t=T)\right]$ at various fixed time slices $t=T$ [^30] NOTE: The Heaviside step function, $\Theta$, used in Eq. (122) has a value of $1/2$ when the argument is zero (i.e. $\Theta(0)=\frac{1}{2}$)
@@ B.1 Analysis of Conditions for EWN
where $z_{i}(x)$ and $t_{i}(x)$ are given in Eq. (12). Along $\chi_{i}$ we have the following relation between $x$ and $|t_{i}(x)-t_{a}|$,
@@
Next we consider the conditions given in Eq. (35). For $\chi_{i}$ we need to consider,
@@
$\begin{split}&\left[|\delta t_{i}|+R_{b}\right]^{2}-\left[b_{i}\sin\theta_{0}-\frac{|\delta t_{i}|}{|k_{i}|}-\frac{b_{1}+b_{2}}{2\sin\theta_{0}}\right]^{2}-z_{i}\left(b_{i}\sin\theta_{0}-\frac{|\delta t_{i}|}{|k_{i}|}\right)^{2}\\ &=\frac{|\delta t_{i}|(a_{i}+b_{i}\sin\theta_{0})}{|k_{i}|}\left[\left[|k_{i}|+\frac{R_{b}}{a_{i}+b_{i}\sin\theta_{0}}\right]^{2}-\frac{\left(2a_{i}+\frac{b_{1}+b_{2}}{\sin\theta_{0}}\right)^{2}}{4(a_{i}+b_{i}\sin\theta_{0})^{2}}\right]<0.\\ \end{split}$ (129)
@@
The set of terms in the second line are zero due to Eq. (133). We arrive at the following conclusion,
@@
So whenever this holds we will know that at least in a neighborhood of the point where the brane and $\chi_{1}$ intersect $\chi_{1}\cap\mathcal{R}_{M}=\varnothing$. [^31] The reader may be concerned that the expression for $k_{1,*}$ is not always going to be real. This is indeed true so the reader should really interpret the inequality is being true when $k_{1,*}\in\mathbb{R}$. In the case when it is not real we can still show that $\chi_{1}\cap\mathcal{R}_{M}=\varnothing$.
@@
The reason for this is because of the result we wrote in Eq. (143), it shows that $-z_{1}(x)|_{k_{1}=0}\geq-z_{1}(x)\Rightarrow z_{\sigma_{L}}(x)-z_{1}(x)|_{k_{1}=0}\geq z_{\sigma_{L}}(x)-z_{1}(x)$. Therefore, satisfying the bounds written in Eq. (147) for $k_{1}=0$ is sufficient to have a bound for general values of $k_{1}$. A simple computation reveals that
@@
    • $|k_{1}|\geq\text{Re}\left[k_{1,*}\right]$. [^34] Here we are combining the cases when $k_{1,*}$ is real or imaginary by simply putting bounds using the real part of $k_{1,*}$.
@@ B.3 Condition for Dominance of Connected Phase
For example, in Figure 16 we plot both $\Delta t_{A,EWN}^{2}-\Delta t_{\text{con}}^{2}$ (left frame) and $\Delta t_{A,EWN}^{2}-\Delta t_{\text{con}}^{2}$ (right frame) as functions of $\theta_{0}$ for fixed $b_{1,2}$ and $a_{+}$, and each colored line in the graph corresponds to a particular choice of $a_{-}$. [^35] The reader might be wondering why we take the difference of squares rather than difference of the quantities themselves. It is just to reduce the number of square roots in our expression which can cause reduce numerical error when we get close to machine precision. We know the quantities themselves are always non-negative so can do this without affecting our conclusions. At first glance, we can see for certain choices of parameters the lines can be above or below the $x$-axis. We can also see that when comparing between the left and right frames, lines of the same color (i.e. the lines that are plotted for same parameters $b_{1,2}$ and $a_{\pm}$) are above or below the $x$-axis for the exact same interval(s) in $\theta_{0}$. This can be seen visually in the plots, and verified through more precise numerical calculations. This is not just true for the particular sample we show but also more generally. This in fact is likely not a coincidence; it is a consequence of another fact that we numerically verified for many different choices of parameters. Which is, the roots of $\Delta t_{A,EWN}-\Delta t_{\text{con}}$ and $\Delta t_{2,EWN}-\Delta t_{\text{con}}$ are precisely given by solutions to $2a_{-}=\ell_{*}$, where we introduced $\ell_{*}$ in Eq. (48). In fact, recalling that by definition $\ell_{*}=2a_{-}\Leftrightarrow\Delta t_{A,EWN}=\Delta t_{2,EWN}$ makes it very clear why the roots match for the same colored lines in Figure 16.